OpenStudy (yacoub1993):

Find the slope of the line through the pair of points. A(2, –3), P(2, 9) A. 0 B. 2/3 C. - 1/3 -1/3 D. undefined

terenzreignz (terenzreignz):

Given any two points \[\large (x_1, y_1) \ and \ (x_2, y_2)\] The slope of the line through them is given by \[\huge \frac{y_2 - y_1}{x_2 - x_1}\]

terenzreignz (terenzreignz):

However, in this case, you seem to end up with \[\frac{9-(-3)}{2-2}=\frac{12}{0}\] Which is undefined. By the way, if the slope of a line is undefined, it means the line is vertical. (if the slope is zero, the line is horizontal)

OpenStudy (yacoub1993):

so i will choose D

terenzreignz (terenzreignz):

That's right.

OpenStudy (yacoub1993):

One more x+y=3 y=2x-15 A. (5 -5) B. (-10, -5) C. (1, 2) D. (6, -3)

OpenStudy (yacoub1993):

i think its "D"

terenzreignz (terenzreignz):

You're right. Again, I hope you arrived at the answer the correct way... (although if you were pressed for time, trial and error may be quicker if you have choices)

OpenStudy (yacoub1993):

-1/3x + y = –1 y = 4 + 1/3x A. infinitely many solutions B. no solution C. no solution D. infinitely many solutions

OpenStudy (yacoub1993):

CHOICE A

OpenStudy (yacoub1993):

CHOICE B

OpenStudy (yacoub1993):

CHOICE C

OpenStudy (yacoub1993):

CHOICE D

OpenStudy (yacoub1993):

Which one do u think

terenzreignz (terenzreignz):

Try to set up both equations so that y goes (alone) on one side, and everything else on the other.

terenzreignz (terenzreignz):

In the second equation, it's already done for you. Let's do the first... y = (1/3)x - 1 Am I right?

terenzreignz (terenzreignz):

Well, anyway, your lines are parallel, and they do not intersect anywhere, ergo, there are no solutions.

OpenStudy (yacoub1993):

so wat would be the answer

terenzreignz (terenzreignz):

No solutions. It's in your choices.

OpenStudy (yacoub1993):

Write a rule for the linear function in the graph. A. y = -4x + 13 B. y = -4x – 13 C. y = 4x – 13 D. y = 1/4x – 13

terenzreignz (terenzreignz):

You have two points marked on your graph. In each of these choices, substitute both points. There is only one equation here that would satisfy both points.

terenzreignz (terenzreignz):

This isn't the proper way to solve this, but time's not on our side... I have to go for now. -- Terence out

OpenStudy (yacoub1993):

what would be the equation if i give up cause i tried it couldnt solve it

terenzreignz (terenzreignz):

Simple, try again... no giving up, lol.

OpenStudy (yacoub1993):

ok bye have a good day